OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y greater than or equal to -3x - 2? (0, -1) (0, 0) (-1, 0) (1, 1)

OpenStudy (anonymous):

An easy way to do this is just to plug the values into the equation and see if it works. (0, -1) \[y \ge -3x - 2\] \[-1 \ge -3(0) - 2\] \[-1\ge-2\] This is true. Just remember that the bigger the numerical value of a negative number is, the small it is. (0,0) \[0 \ge -3(0) - 2\] \[0\ge-2\] Again, this is true. (-1, 0) \[0 \ge -3(-1) - 2\] \[0\ge1\] This is not true, so this is your answer

OpenStudy (anonymous):

ty